# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

### v1

@@ A.1 Proof of Theorem 1
Now, we have that $A(z)$ is identically equal to zero thanks to our condition. We claim that the partial integral that depends of $B(z)$ and $D(z)$ is finite if $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$. In fact, we have, for $z\to\infty$,
$|B(z)|$ $\leq c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}+c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\leq c_{6}\min\{z^{\lambda_{1}+2\delta},z^{\lambda_{0}+2\delta}\}$
and
$|D(z)|$ $\leq\mathcal{O}(z^{-1})\{c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}+c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\}$
$\leq c_{7}\min\{z^{\lambda_{1}+2\delta},z^{\lambda_{0}+2\delta}\},$
with $c_{i}$ being positive and finite constants, $i=6,7$. Thus, if $\frac{\sigma_{0}^{2}L_{0}}{a_{0}^{2\delta-d}}=\frac{\sigma_{1}^{2}L_{1}}{a_{1}^{2\delta-d}}$, we have
$\int_{c}^{\infty}$ $z^{d-1}\left(\frac{\widehat{\mathcal{GH}}_{\delta,\beta_{1},\gamma_{1},a_{1}}(z)-\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}{\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}\right)^{2}{\rm d}z$ (33)
$\leq\int_{c}^{\infty}z^{d-1}(B(z)+D(z))^{2}{\rm d}z$
$\leq\int_{c}^{\infty}z^{d-1}\Big{(}B(z)^{2}+D(z)^{2}+2B(z)D(z)\Big{)}{\rm d}z.$
We conclude that (33) is finite if $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$. The sufficient part of our claim is thus proved.
The necessary part follows the proof of Zhang, (2004). We suppose that $\frac{d+1}{2}+3\delta\leq\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})$, $\frac{\sigma_{0}^{2}L_{0}}{a_{0}^{2\delta-d}}\neq\frac{\sigma_{1}^{2}L_{1}}{a_{1}^{2\delta-d}}$ and let $\sigma_{2}^{2}=\sigma_{0}^{2}L_{0}\left(\frac{a_{1}}{a_{0}}\right)^{2\delta-d}$. Then $\mathcal{GH}_{\delta,\beta_{0},\gamma_{0},a_{0}}$ and $\mathcal{GH}_{\delta,\beta_{1},\gamma_{1},a_{2}}$ induce two equivalent Gaussian measures. This, we need to show that $\mathcal{GH}_{\delta,\beta_{1},\gamma_{1},a_{2}}$ and $\mathcal{GH}_{\delta,\beta_{1},\gamma_{1},a_{1}}$ define two orthogonal Gaussian measures. The rest of the proof follows the same arguments in Zhang, (2004). ∎
@@ A.2 Proof of Theorem 2
Let us consider the $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ model and, for the moment, let us assume that $l$ is a real number. The sufficient conditions (a), (b), (c) given in Section 3, under the parameterization (15) are:
@@
Solving the previous system of inequalities, we derive the following sufficient conditions:
  1. 1.
    For $d=1$
    • a)
      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l\leq-\frac{(1+2\kappa)}{2}$
    • b)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(1+2\kappa)}{2}<l\leq\frac{(1+2\kappa)}{2}$.
    • c)
      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l>\frac{(1+2\kappa)}{2}$.
  2. 2.
    For $d>1$
    • a)
      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l<-\frac{(d+2\kappa)}{2}$
    • b)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(d+2\kappa)}{2}\leq l\leq\frac{(d+2\kappa)}{2}$.
    • c)
      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l>\frac{(d+2\kappa)}{2}$.
Necessary conditions can be obtained following Cho and Yun, (2018, Theorem 6.1), who show that if $\mathstrut{}_{1}F_{2}\left(a;b;c;-\frac{x^{2}}{4}\right)\geq 0$, then $b>a$, $c>a$, $b+c\geq 3a+\frac{1}{2}$. Using the spectral density (11), under the parameterization (15), we have
  1. (a)
    $\mu>0$;
  2. (b)
    $\mu>-2l$;
  3. (c)
    $\mu+l\geq\frac{d}{2}+\kappa+1$.
Solving the system of inequalities (under the restriction $\kappa>-\frac{1}{2}$) for $d\geq 1$, we derive the following necessary conditions:
  1. 1.
    For $d=1$
    • a)
      $\mu>-2l$ and $l\leq-\frac{(3+2\kappa)}{2}$
    • b)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(3+2\kappa)}{2}<l\leq\frac{(3+2\kappa)}{2}$.
    • d)
      $\mu>0$ and $l>\frac{(3+2\kappa)}{2}$.
  2. 2.
    For $d>1$
    • a)
      $\mu>-2l$ and $l<-\frac{(d+2\kappa+2)}{2}$
    • c)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(d+2\kappa+2)}{2}\leq l<\frac{(d+2\kappa+2)}{2}$.
    • d)
      $\mu>0$ and $l\geq\frac{(d+2\kappa+2)}{2}$.
Putting together the necessary and sufficient conditions for the cases $d=1$ and $d>1$, we obtain the following conditions:
  1. 1.
    For $d=1$
    • a)
      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l\leq-\frac{(1+2\kappa)}{2}$
    • b)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(1+2\kappa)}{2}<l\leq\frac{(1+2\kappa)}{2}$.
    • c)
      $\mu\geq\sqrt{2\kappa+l^{2}+2}-l$ and $l>\frac{(1+2\kappa)}{2}$.
  2. 2.
    For $d>1$
    • a)
      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l<-\frac{(d+2\kappa)}{2}$
    • b)
      $\mu\geq\delta-(l-\frac{1}{2})$ and $-\frac{(d+2\kappa)}{2}\leq l\leq\frac{(d+2\kappa)}{2}$.
    • c)
      $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l>\frac{(d+2\kappa)}{2}$.
We observe that the necessary and sufficient conditions are symmetric around $l=0$, since $\mathcal{GH}_{\delta,\beta,\gamma,a}(x)=\mathcal{GH}_{\delta,\gamma,\beta,a}(x)$. Specifically, for $d=1$, conditions a) and c) are symmetric when $l<-\frac{(1+2\kappa)}{2}$ and $l>\frac{(1+2\kappa)}{2}$, respectively, while condition b) is symmetric when $-\frac{(d+2\kappa)}{2}<l\leq 0$ and $0\leq l\leq\frac{(d+2\kappa)}{2}$. These same symmetry patterns are also observed for the case $d>1$.
Using the symmetry of the conditions, we put together the necessary and sufficient conditions only for $l\geq 0$ obtaining the main result, that is for $\kappa>-\frac{1}{2}$ the model is valid if and only if
  1. 1.
    $\mu\geq\delta-(l-\frac{1}{2})$ and $0\leq l\leq\frac{(d+2\kappa)}{2}$
  2. 2.
    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ and $l>\frac{(d+2\kappa)}{2}$.
@@ A.3 Proof of Theorem 3
The result stems from formula 2.21.1.11 of Prudnikov et al., (1990) and the expressions (12) and (16) of the hypergeometric correlation and of the integral range.
@@ A.4 Proof of Theorem 4
Hereafter, without loss of generality consider a unit compact support $a=1$. Let us define $LA_{\kappa,d}(\mu,l):=log(A_{\kappa,\mu,l,d,1})$ where
$A_{\kappa,\mu,l,d,1}=\frac{\pi^{d/2}\Gamma((d+1)/2+\kappa)\Gamma\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\Gamma(\kappa+0.5)\Gamma\left(\frac{\mu+1}{2}+\kappa+d/2\right)\Gamma\left(\frac{\mu+1}{2}+\kappa+l+d/2\right)}.$
is the integral range.
[…]
@@
    $\mu\geq\delta-l+0.5\quad$ if $\quad 0\leq l\leq d/2+\kappa$
@@
    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l\quad$ if $\quad l>d/2+\kappa$.
This is a maximization problem over a bounded interval. Hereafter, with $\psi(x)=\frac{\text{d}log\Gamma(x)}{\text{d}x}$ we denote the digamma function which is strictly increasing for $x>0$. The fist derivative of $LA_{\kappa,d}(\mu,l)$ with respect to $\mu$ is given by
$\frac{\text{d}LA_{\kappa,d}(\mu,l)}{\text{d}\mu}=0.5[\psi(x)-\psi(x+d/2)+\psi(x+l)-\psi(x+l+d/2)],$
where $x=\frac{\mu+1}{2}+\kappa$. It is a negative function since $\psi(\cdot)$ is strictly increasing and this implies that the function $LA_{\kappa,d}(\mu,l)$ is monotone non-increasing as a function of $\mu$. As a consequence the maximum is achieved by $LA_{\kappa,d}(\mu,l)$ on the lower boundary of region 1) and region 2), that is
  1. 1.
[…]
  2. 2.
[…]
where $\mu_{1}(l)=\delta-l+0.5$ and $\mu_{2}(l)=\sqrt{2\kappa+l^{2}+d+1}-l$. This proves Theorem 4.
@@ A.5 Proof of Theorem 5
We focus on the functions $LA_{\kappa,d}(\mu_{i}(l),l)$ for $i=1,2$ defined on the proof of Theorem 4 and the goal is to find the maximum with respect to $l$.
If $i=1$, the derivative of $LA_{\kappa,d}(\mu_{1}(l),l)$ for $0\leq l\leq d/2+\kappa$ is given by
$\frac{\text{d}LA_{\kappa,d}(\mu_{1}(l),l)}{\text{d}l}=0.5[\psi(x+l/2)-\psi(x-l/2)+\psi(x-l/2+d/2)-\psi(x+l/2+d/2)],$
where $x=\frac{d}{4}+\frac{3\kappa}{2}+1$. It can be shown that this function is non-negative using the fact that the derivative of the digamma function $\frac{\text{d}\psi(x)}{\text{d}x}$ is a decreasing function. This implies that the maximum of $LA_{\kappa,d}(\mu_{1}(l),l)$ is achieved on the upper bound of the parametric space, that is $l=d/2+\kappa$.
Using similar arguments, if $i=2$ it can be shown that the derivative of $LA_{\kappa,d}(\mu_{2}(l),l)$ for $l>d/2+\kappa$ is non-positive. As a consequence the maximum is achieved on the lower bound of the parametric space, that is when $l=d/2+\kappa+\epsilon$ where $\epsilon$ an arbitrary small positive constant.
Finally we can prove that $LA_{\kappa,d}(\mu_{1}(d/2+\kappa),d/2+\kappa))=LA_{\kappa,d}(1,d/2+\kappa)$ is greater than $LA_{\kappa,d}(\mu_{2}(d/2+\kappa+\epsilon),d/2+\kappa+\epsilon))=LA_{\kappa,d}(1-g(\epsilon),d/2+\kappa+\epsilon)$ where $g(\epsilon)$ is a positive constant.

### v2

@@ A.1 Proof of Theorem 1
Assume now (10), i.e.
$\frac{\sigma_{0}^{2}L(\beta_{0},\gamma_{0})}{a_{0}^{2\delta-d}}=\frac{\sigma_{1}^{2}L(\beta_{1},\gamma_{1})}{a_{1}^{2\delta-d}}.$
This is equivalent to
$c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}=c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta}K_{0},$
so that $A(z)\equiv 0$.
Using $(x+y)^{2}\leq 2x^{2}+2y^{2}$, it suffices to check
$\int_{c}^{\infty}z^{d-1+2(\lambda_{i}+2\delta)}\,dz<\infty,\qquad i=0,1,$
which holds iff
$\beta_{i}+\gamma_{i}>\frac{d+1}{2}+3\delta.$
Therefore, under
$\min(\beta_{1}+\gamma_{1},\beta_{0}+\gamma_{0})>\frac{d+1}{2}+3\delta,$
condition (26) holds and the sufficient part is proved.
The necessary part follows from Zhang (2004).
@@ A.2 Proof of Theorem 2
Consider the model $\mathcal{GH}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}$ in dimension $d$, where $\delta=\frac{d+1}{2}+\kappa$, $\kappa>-\frac{1}{2}$, $l\geq 0$.
Sufficiency. Under $\beta=\delta+\frac{\mu}{2}$, $\gamma=\delta+\frac{\mu}{2}+l$, the sufficient conditions of Section 3 become:
@@
Condition (b) is equivalent to the quadratic inequality
$\mu^{2}+2l\mu-(d+2\kappa+1)\geq 0,$
whose positive branch gives
$\mu\geq-l+\sqrt{l^{2}+d+2\kappa+1}.$
Only the positive branch is admissible since condition (c) implies $\mu>-l$ when $l\geq 0$. For each fixed $l$, the sufficient conditions therefore reduce to
$\mu\geq\max\!\Big\{\frac{d}{2}+\kappa+1-l,\;-l+\sqrt{l^{2}+d+2\kappa+1}\Big\}.$
A direct algebraic comparison shows that
$\frac{d}{2}+\kappa+1-l\geq-l+\sqrt{l^{2}+d+2\kappa+1}\quad\Longleftrightarrow\quad 0\leq l\leq\frac{d}{2}+\kappa.$
Hence condition (c) is binding when $0\leq l\leq\frac{d}{2}+\kappa$, giving $\mu\geq\delta-l+\frac{1}{2}$, and condition (b) is binding when $l>\frac{d}{2}+\kappa$, giving $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$. This yields exactly the two regimes in points 1 and 2.
Necessity. Assume $\mathcal{GH}_{\delta,\beta,\gamma,a}\in\Phi_{d}$. Since the Gamma prefactors in the spectral density are positive whenever $\beta>\delta$ and $\gamma>\delta$, nonnegativity of the spectral density requires
$_{1}F_{2}\!\left(\delta;\beta,\gamma;-\frac{(za)^{2}}{4}\right)\geq 0\quad\text{for all }z\geq 0.$
By Cho and Yun (2018, Theorem 3.1), this implies $\beta>\delta$, $\gamma>\delta$, and $\beta+\gamma\geq 3\delta+\frac{1}{2}$. The third condition, under the parameterization $\beta=\delta+\frac{\mu}{2}$, $\gamma=\delta+\frac{\mu}{2}+l$, gives
$\mu+l\geq\frac{d}{2}+\kappa+1,\quad\text{i.e.,}\quad\mu\geq\delta-l+\frac{1}{2},$
which is the necessary lower bound for $\mu$ in both regimes. In Regime 1 ($0\leq l\leq\frac{d}{2}+\kappa$), this coincides with the sufficient condition in point 1, so the result is an if and only if. In Regime 2 ($l>\frac{d}{2}+\kappa$), this provides a necessary lower bound that is strictly weaker than the sufficient condition in point 2; whether the gap can be closed remains an open question.
Reduction to $l\geq 0$. The model $\mathcal{GH}_{\delta,\beta,\gamma,a}$ is invariant under interchanging $\beta$ and $\gamma$, which corresponds to $l\mapsto-l$. The admissible region is therefore symmetric with respect to $l=0$, and restricting to $l\geq 0$ entails no loss of generality.
This completes the proof. ∎
@@ A.3 Proof of Theorem 3
Without loss of generality set $a=1$; the general case follows by the substitution $x\mapsto ax$, which contributes the factor $a^{d}$. Using the change of variable $t=1-x^{2}$ (so $x^{d-1}\,dx=\tfrac{1}{2}(1-t)^{(d-2)/2}\,dt$) in the definition of the integral range (12), and substituting (8), one obtains
$A_{\kappa,\mu,l,1,d}$ $=\frac{\pi^{d/2}C_{0}}{2^{d}\Gamma(d/2)}\int_{0}^{1}(1-t)^{\frac{d-2}{2}}\,t^{\mu+\frac{d-1}{2}+2\kappa}$ (28)
$\quad\times{}_{2}F_{1}\!\left(\frac{\mu}{2},\frac{\mu+d}{2}+\kappa;\mu+\frac{d+1}{2}+2\kappa;t\right)\,dt.$
where $C_{0}=\Gamma\!\left(\kappa+\frac{\mu+1}{2}\right)\Gamma\!\left(2\kappa+\frac{d+\mu+1}{2}\right)\big/\!\left[\Gamma\!\left(\mu+\frac{d+1}{2}+2\kappa\right)\Gamma\!\left(\kappa+\tfrac{1}{2}\right)\right]$.
We apply formula 2.21.1.11 of Prudnikov et al. (1990),
$\int_{0}^{1}t^{\sigma-1}(1-t)^{\rho-1}\,{}_{2}F_{1}(a,b;c;t)\,dt=\frac{\Gamma(\rho)\Gamma(\sigma)}{\Gamma(\rho+\sigma)}\,{}_{3}F_{2}(a,b,\sigma;\,c,\,\rho+\sigma;\,1),$ (29)
with the identifications
$\sigma=\mu+\tfrac{d+1}{2}+2\kappa,\quad\rho=\tfrac{d}{2},\quad a=\tfrac{\mu}{2},\quad b=\tfrac{\mu+d}{2}+\kappa,\quad c=\mu+\tfrac{d+1}{2}+2\kappa.$
Since $c=\sigma$, the third upper parameter and the first lower parameter of $_{3}F_{2}$ coincide, so the series telescopes term by term:
$_{3}F_{2}(a,b,\sigma;\,\sigma,\,\rho+\sigma;\,1)=\sum_{k=0}^{\infty}\frac{(a)_{k}(b)_{k}(\sigma)_{k}}{k!\,(\sigma)_{k}(\rho+\sigma)_{k}}=\sum_{k=0}^{\infty}\frac{(a)_{k}(b)_{k}}{k!\,(\rho+\sigma)_{k}}={}_{2}F_{1}(a,b;\,\rho+\sigma;\,1).$
Gauss’s summation theorem then gives, provided $\rho+\sigma-a-b>0$,
$_{2}F_{1}(a,b;\rho+\sigma;1)=\frac{\Gamma(\rho+\sigma)\,\Gamma(\rho+\sigma-a-b)}{\Gamma(\rho+\sigma-a)\,\Gamma(\rho+\sigma-b)}.$
One verifies directly that
$\rho+\sigma-a-b=\kappa+\tfrac{1}{2}+l>0\quad(\kappa>-\tfrac{1}{2},\;l\geq 0),$
$\rho+\sigma-a=\tfrac{\mu+1}{2}+\kappa+d,\qquad\rho+\sigma-b=\tfrac{\mu+1}{2}+\kappa+\tfrac{d}{2}.$
Substituting back into (28), the factors $\Gamma(\sigma)$ and $\Gamma(\rho+\sigma)$ cancel with those arising from (29) and from $C_{0}$. Using the duplication formula $\Gamma(z)\Gamma(z+\tfrac{1}{2})=\sqrt{\pi}\,2^{1-2z}\Gamma(2z)$ to simplify the remaining Gamma factors, one obtains
$A_{\kappa,\mu,l,1,d}=\frac{\pi^{d/2}\,\Gamma\!\left(\frac{d+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\,\Gamma\!\left(\kappa+\frac{1}{2}\right)\Gamma\!\left(\frac{\mu+1+d}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+1+d}{2}+\kappa+l\right)},$
which, after restoring the factor $a^{d}$, gives (14). ∎
@@ A.4 Proof of Theorem 4
Proof.
Without loss of generality, set the compact support to $a=1$. Define
$LA_{\kappa,d}(\mu,l):=\log(A_{\kappa,\mu,l,d,1}),$
where
$A_{\kappa,\mu,l,d,1}=\frac{\pi^{d/2}\Gamma((d+1)/2+\kappa)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l\right)}{2^{d}\Gamma(\kappa+0.5)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+\frac{d}{2}\right)\,\Gamma\!\left(\frac{\mu+1}{2}+\kappa+l+\frac{d}{2}\right)}.$
We maximize $LA_{\kappa,d}(\mu,l)$ over the admissible set in $\mu$ for a fixed $l\geq 0$, as given by Theorem 2. Let $\delta=\frac{d}{2}+\kappa$. Then
@@
    $\mu\geq\delta-l+\tfrac{1}{2}$ if $0\leq l\leq\delta$;
@@
    $\mu\geq\sqrt{2\kappa+l^{2}+d+1}-l$ if $l>\delta$.
This is a maximization problem over the half-line $\mu_{\min}(l),\infty)$.
Let $\psi(x)=\frac{d}{dx}\log\Gamma(x)$ be the digamma function, which is strictly increasing for $x>0$. Set $x=\frac{\mu+1}{2}+\kappa$. Under the validity conditions we have $x>0$ and $x+l>0$. Differentiating with respect to $\mu$ yields
$\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)=\frac{1}{2}\Big[\psi(x)-\psi(x+\tfrac{d}{2})+\psi(x+l)-\psi(x+l+\tfrac{d}{2})\Big].$
Since $\psi$ is increasing and $\tfrac{d}{2}>0$, we have $\psi(x)-\psi(x+\tfrac{d}{2})\leq 0$ and $\psi(x+l)-\psi(x+l+\tfrac{d}{2})\leq 0$, hence $\frac{d}{d\mu}LA_{\kappa,d}(\mu,l)\leq 0$ for all admissible $\mu$. Therefore $LA_{\kappa,d}(\mu,l)$ is non-increasing in $\mu$, and its maximum is achieved at the smallest admissible value of $\mu$, i.e.
$\mu_{1}(l)=\delta-l+\tfrac{1}{2}\quad(0\leq l\leq\delta),\qquad\mu_{2}(l)=\sqrt{2\kappa+l^{2}+d+1}-l\quad(l>\delta).$
This proves Theorem 4. ∎
@@ A.5 Proof of Theorem 5
Proof.
Let $\delta=\frac{d}{2}+\kappa$. Since $\log$ is strictly increasing, maximizing $A_{\kappa,\mu,l,a,d}$ is equivalent to maximizing
$LA_{\kappa,d}(\mu,l):=\log A_{\kappa,\mu,l,d,1},$
and the factor $a^{d}$ plays no role for maximizers.
By Theorem 4, for each fixed $l$ the maximizer in $\mu$ is $\mu=\mu_{1}(l)$ for $0\leq l\leq\delta$ and $\mu=\mu_{2}(l)$ for $l>\delta$. Hence it suffices to study the two one–variable functions
$F_{1}(l):=LA_{\kappa,d}(\mu_{1}(l),l)\quad\text{on }[0,\delta],\qquad F_{2}(l):=LA_{\kappa,d}(\mu_{2}(l),l)\quad\text{on }(\delta,\infty).$
Write $\psi=\Gamma^{\prime}/\Gamma$ for the digamma function. Recall that $\psi$ is concave on $(0,\infty)$, so for every $h>0$ the symmetric increment
$\Delta_{h}(t):=\psi(t+h)-\psi(t-h)$
is decreasing in $t$ (because $\Delta_{h}^{\prime}(t)=\psi^{\prime}(t+h)-\psi^{\prime}(t-h)\leq 0$).
Case $i=1$ ($0\leq l\leq\delta$).
For $0\leq l\leq\delta$, $\mu_{1}(l)=\delta-l+\tfrac{1}{2}$. Set
$u(l):=\frac{\mu_{1}(l)+1}{2}+\kappa=\frac{d}{4}+\frac{3\kappa}{2}+\frac{3}{4}-\frac{l}{2}.$
Using the explicit form of $A_{\kappa,\mu,l,d,1}$, $F_{1}(l)$ differs from
$\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\Big(u(l)+\frac{d}{2}\Big)-\log\Gamma\Big(u(l)+l+\frac{d}{2}\Big)$
only by an additive constant independent of $l$. Differentiating gives
$F_{1}^{\prime}(l)$ $=u^{\prime}(l)\Big[\psi(u(l))+\psi(u(l)+l)-\psi\Big(u(l)+\frac{d}{2}\Big)-\psi\Big(u(l)+l+\frac{d}{2}\Big)\Big]$
$\qquad+\psi(u(l)+l)-\psi\Big(u(l)+l+\frac{d}{2}\Big),$
where $u^{\prime}(l)=-\tfrac{1}{2}$. A simple rearrangement yields
$F_{1}^{\prime}(l)=\frac{1}{2}\Big[\psi(u(l)+l)-\psi(u(l))+\psi\Big(u(l)+\frac{d}{2}\Big)-\psi\Big(u(l)+l+\frac{d}{2}\Big)\Big].$
Now set
$x:=\frac{d}{4}+\frac{3\kappa}{2}+\frac{3}{4}\quad\Longrightarrow\quad u(l)=x-\frac{l}{2},\qquad u(l)+l=x+\frac{l}{2}.$
Therefore
$F_{1}^{\prime}(l)=\frac{1}{2}\Big[\psi(x+\tfrac{l}{2})-\psi(x-\tfrac{l}{2})+\psi(x-\tfrac{l}{2}+\tfrac{d}{2})-\psi(x+\tfrac{l}{2}+\tfrac{d}{2})\Big]=\frac{1}{2}\Big(\Delta_{l/2}(x)-\Delta_{l/2}(x+\tfrac{d}{2})\Big).$
Since $\Delta_{l/2}$ is decreasing and $x+\tfrac{d}{2}>x$, we have $\Delta_{l/2}(x)\geq\Delta_{l/2}(x+\tfrac{d}{2})$, hence $F_{1}^{\prime}(l)\geq 0$ on $0,\delta]$. Thus $F_{1}$ is non-decreasing on $0,\delta]$ and its maximum is attained at $l=\delta$.
Case $i=2$ ($l>\delta$).
For $l>\delta$, $\mu_{2}(l)=\sqrt{l^{2}+(2\kappa+d+1)}-l$. Define
$u(l):=\frac{\mu_{2}(l)+1}{2}+\kappa>0,\qquad x(l):=u(l)+\frac{l}{2}=\kappa+\frac{1+\sqrt{l^{2}+(2\kappa+d+1)}}{2}.$
Then $u(l)=x(l)-\tfrac{l}{2}$ and $u(l)+l=x(l)+\tfrac{l}{2}$. As above, up to an additive constant independent of $l$,
$F_{2}(l)=\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\Big(u(l)+\frac{d}{2}\Big)-\log\Gamma\Big(u(l)+l+\frac{d}{2}\Big).$
Differentiating (same computation as in Case $i=1$ but with general $u$) gives the identity
$F_{2}^{\prime}(l)=\Big(x^{\prime}(l)\Big)\Big(\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\Big)-\frac{1}{2}\Big(\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\Big),$
i.e.
$F_{2}^{\prime}(l)=\Big(x^{\prime}(l)-\frac{1}{2}\Big)\Big(\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\Big).$
Now observe that
$x^{\prime}(l)=\frac{1}{2}\cdot\frac{l}{\sqrt{l^{2}+(2\kappa+d+1)}}\in\Big(0,\frac{1}{2}\Big),$
so $x^{\prime}(l)-\frac{1}{2}<0$ for all $l>\delta$. On the other hand, as in Case $i=1$, concavity of $\psi$ implies $\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\geq 0$. Therefore $F_{2}^{\prime}(l)\leq 0$ for all $l>\delta$. Hence $F_{2}$ is non-increasing on $(\delta,\infty)$ and
$\sup_{l>\delta}LA_{\kappa,d}(\mu_{2}(l),l)=\lim_{l\downarrow\delta}LA_{\kappa,d}(\mu_{2}(l),l),$
equivalently
$\sup_{l>\delta}A_{\kappa,\mu_{2}(l),l,a,d}=\lim_{l\downarrow\delta}A_{\kappa,\mu_{2}(l),l,a,d}.$
Finally, $\mu_{2}(l)\to 1$ as $l\downarrow\delta$ and $\mu_{1}(\delta)=1$. By continuity of $LA_{\kappa,d}(\mu,l)$ on the admissible region,
$\lim_{l\downarrow\delta}LA_{\kappa,d}(\mu_{2}(l),l)=LA_{\kappa,d}(1,\delta)=LA_{\kappa,d}(\mu_{1}(\delta),\delta).$
Therefore the maximum of $A_{\kappa,\mu_{1}(l),l,a,d}$ over $0\leq l\leq\delta$ is attained at $l=\delta$, and the supremum of $A_{\kappa,\mu_{2}(l),l,a,d}$ over $l>\delta$ equals the same value and is approached as $l\downarrow\delta$. Consequently, the global maximum over the admissible ranges is achieved at $(\mu,l)=(1,\delta)$, while in the regime $l>\delta$ it is not attained but only approached in the limit $l\downarrow\delta$. This proves Theorem 5. ∎